determine the period of the function y= 2cos 3x and select the correct option
a. pi/3
b. 2pi/3
c. pi
d. 4pi/3

Answer by stanbon(75887)

determine the period of the function y= 2cos 3x and select the correct option
======
Form:: y = a*cos(bx)
---------------------------
Period = (2pi)/b = (2pi)/3 = (2pi/3)
----------------
Cheers,
Stan H.
---------------
a. pi/3
b. 2pi/3
c. pi
d. 4pi/3